# Mod $\ell$ non-vanishing of self-dual Hecke $L$-values over CM fields and applications

### v1

@@ 4.4.1. Inert case
where $\doteq$ means up to an $\ell$-adic unit. Hence, in view of the explicit Waldspurger formula as in  Theorem 3.12, Proposition 3.14 and Lemma 4.13, it suffices to show that
@@
for all but finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$.
For $n\gg 0$, we have
${\rm G}_{n}=(D_{1}\times D_{0})\cdot\Gamma_{{\mathfrak{p}},n},$
where $D_{1}$, $D_{0}$ are as in §4.1.1 and $\Gamma_{{\mathfrak{p}},n}:=\Gamma_{{\mathfrak{p}}}/(1+{\mathfrak{p}}^{n}O_{K_{\mathfrak{p}}})$. Note that $|D_{0}|$ is an $\ell$-adic unit by our assumption on $\ell$. It follows from the definition of $f$ that
$\sum_{d\in D_{0}}\pi(\iota(d))H_{\ell}=|D_{0}|\cdot H_{\ell}.$ (4.4)
Let $p$ be the rational prime below $\mathfrak{p}$}. Let $\nu:\Gamma_{{\mathfrak{p}},n}\to\mu_{p^{\infty}}$ be a character of conductor ${\mathfrak{p}}^{n}$ with $n>a$. Put
$C_{n}=\Gamma_{{\mathfrak{p}},n}[p^{a}].$
Note that $C_{n}={\mathrm{Ker}}({\rm G}_{n}\to{\rm G}_{n-a})$. Let ${\bf k}_{\ell}(\nu)$ be the field generated by the values of $\nu$ over $\bf k$$_{\ell}$. Since $\bf k$$_{\ell}$ contains $\mu_{p}$, $d_{\nu}:=[{\bf k}_{\ell}(\nu):{\bf k}_{\ell}]$ is a $p$-power, and for a $p$-power root of unity $\zeta\in{\bf k}_{\ell}(\nu)$, we have
In the following, we identify $F_{{\mathfrak{p}}}$ with ${\mathbb{Q}}_{p}$ and the matrices appearing in the Hecke action are seen as elements of $B_{{\mathfrak{p}}}={\mathrm{GL}}_{2}({\mathbb{Q}}_{p})$.
In view of the above discussion, it follows that for $n\geq\max\{2m,2a\}$ and $s\in\widehat{K}^{\times}$, we have
${\rm Tr}_{{\bf k}_{\ell}(\nu)/{\bf k}_{\ell}}(\nu\chi_{0}(s^{-1})\cdot\nu(\Theta_{\chi_{0},n}(\pi_{\lambda}))\pmod{\varpi})$= $|D_{0}|d_{\nu}\cdot\sum_{\tau\in D_{1}}{\chi_{0}}(\tau)\sum_{y\in{\mathbb{Z}}/p^{a}\mathbb{Z}}\zeta_{\nu}^{y}H_{\ell}\left(x_{{\mathfrak{c}},n}(s\tau)\begin{pmatrix}1&\frac{y}{p^{a}}\\ 0&1\\ \end{pmatrix}\right)$ (4.6)
for a primitive $p^{a}$-th root of unity $\zeta_{\nu}$.
Define $\widetilde{H}_{\ell}\in{\rm M}_{2}({\bf k}_{\ell}(\nu))$ by
$\widetilde{H}_{\ell}(g):=\sum_{y\in{\mathbb{Z}}/p^{a}{\mathbb{Z}}}\zeta_{\nu}^{y}\rho\left(\begin{pmatrix}1&\frac{y}{p^{a}}\\ 0&1\\ \end{pmatrix}\right)H_{\ell}(g).$ (4.7)
Then $\widetilde{H}_{\ell}\in{\rm M}_{2}(\widetilde{U}^{\prime},{\bf k}_{\ell}(\nu))$ for $\widetilde{U}^{\prime}:=\{g\in U^{\prime}\ |\ g_{p}\equiv\begin{pmatrix}1&*\\ 0&1\end{pmatrix}\pmod{p^{2a}}\}$. By (4.6), we have
${\rm Tr}_{{\bf k}_{\ell}(\nu)/{\bf k}_{\ell}}(\nu\chi_{0}(s^{-1})\cdot\nu(\Theta_{\chi_{0},n}(\pi_{\lambda}))\pmod{\varpi})=|D_{0}|d_{\nu}\cdot\sum_{\tau\in D_{1}}{\chi_{0}}(\tau)\widetilde{H}_{\ell}(x_{{\mathfrak{c}},n}(s\tau)).$ (4.8)
We next show that $\widetilde{H}_{\ell}$ is non-Eisenstein.
To begin, (4.3) holds by Lemma 3.11. Therefore, since $\pi_{\lambda}$ is CM and $(\ell,2D_{K/F})=1$, the non-vanishing of $H_{\ell}$ implies that it is non-Eisenstein by Lemma 4.9. A simple calculation shows that
$\sum_{t\in(\mathbb{Z}/p^{a}\mathbb{Z})^{\times}}\rho\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\widetilde{H}_{\ell}$= $p^{a}H_{\ell}-{p^{a-1}}\rho\left(\begin{pmatrix}p^{-1}&\\ &1\end{pmatrix}\right)U_{p}H_{\ell}.$ (4.9)
$\sum_{t\in(\mathbb{Z}/p^{a}\mathbb{Z})^{\times}}\rho\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\widetilde{H}_{\ell}=p^{a}H_{\ell}-{p^{a-1}}\rho\left(\begin{pmatrix}p^{-1}&\\ &1\end{pmatrix}\right)T_{\mathfrak{p}}H_{\ell}+{p^{a-1}}\rho\left(\begin{pmatrix}p^{-2}&\\ &1\end{pmatrix}\right)H_{\ell}$ (4.10)
Since $H_{\ell}$ is either a $U_{\mathfrak{p}}$-eigenform or a $T_{\mathfrak{p}}$-eigenform with an $\ell$-integral eigenvalue, it follows that $\widetilde{H}_{\ell}$ is non-Eisenstein by Lemma 4.6. Then we may suppose that $\widetilde{H}_{\ell}^{\varepsilon}$ is non-Eisenstein for some $\varepsilon\in\{+,-\}$.
In view of (4.8) and Proposition 4.4, it thus follows that
${v}_{\ell}(\nu(\Theta_{\chi_{0},n}(\pi_{\lambda})))=0$
for all but finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{{\varepsilon_{0}}}$, where $\varepsilon_{0}$} denotes the sign of $(-1)^{n}\varepsilon(\lambda\chi_{0})$ so that
$(-1)^{n}\eta_{K/F}({\mathfrak{c}})\equiv\varepsilon\pmod{2}$
by Proposition 4.4 (cf. Lemma 4.11).
@@
for all but finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{-}$. But these central $L$-values vanish since $\varepsilon(\lambda\chi_{0}\nu)=-1$ for any $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{-}$. Therefore $\varepsilon_{0}=+$ and the parity $n$ in Proposition 4.4 satisfies $(-1)^{n}\varepsilon(\lambda\chi_{0})=+1.$ Moreover, it follows from Lemma 4.11 that
$\varepsilon(\lambda\chi_{0})=\varepsilon(\lambda)\eta_{K/F}\circ{\rm{N}}(\prod_{v|{\mathfrak{c}}{\mathfrak{p}}}\varsigma_{v}^{({\mathrm{ord}}_{v}{\mathfrak{c}})}).$
Hence, $\widetilde{H}_{\ell}^{\varepsilon(\lambda)}$ is non-Eisenstein and $\widetilde{H}_{\ell}^{-\varepsilon(\lambda)}$ is Eisenstein.

### v2

@@ 4.4.1. Inert case
and in view of the explicit Waldspurger formula as in  Theorem 3.14, Proposition 3.16 and Lemma 4.14, it suffices to show that
${v}_{\ell}(\nu(\Theta_{\chi_{0},n}(\pi_{\lambda})))=0$ (4.8)
for all but finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$. In the following we first reduce this non-vanishing to that of the right hand side of  (4.10) by considering the Galois average of toric periods.
Galois average. Let $\nu\in\Xi_{\mathfrak{p}}$ be a character of conductor ${\mathfrak{p}}^{n}$ with $n>m_{\lambda}$. Let ${\bf k}_{\ell}(\nu)\subset\overline{{\mathbb{F}}}_{\ell}$ be the subfield generated by the values of $\nu$ over $\bf k$$_{\ell}$. Since $\bf k$$_{\ell}$ contains $p$-th roots of unity, $d_{\nu}:=[{\bf k}_{\ell}(\nu):{\bf k}_{\ell}]$ is a $p$-power. For a $p$-power root of unity $\zeta\in{\bf k}_{\ell}(\nu)$, we have
Let ${\rm{G}}_{{\mathfrak{c}},n}^{\prime}$ be the subgroup of ${\rm G}_{{\mathfrak{c}},n}$ given by
${\rm G}_{{\mathfrak{c}},n}^{\prime}=\{t\in{\rm G}_{{\mathfrak{c}},n}\ |\ \nu(t)\in\mu_{p^{s}}\}.$
In view of (4.9) and (4.7), for $a\in\widehat{K}^{\times}$, we have
${\rm Tr}_{{\bf k}_{\ell}(\nu)/{\bf k}_{\ell}}(\nu\chi_{0}(a^{-1})\cdot\nu(\Theta_{\chi_{0},n}(\pi_{\lambda}))\pmod{\varpi})=d_{\nu}\cdot\sum_{t\in{\rm G}_{{\mathfrak{c}},n}^{\prime}}\chi_{0}\nu(t)\langle{x,\widehat{\varphi}(x_{{\mathfrak{c}},n}(at))}\rangle_{k-2\Sigma}\pmod{\varpi}$. (4.10)
Hence the non-vanishing (4.8) is a consequence of that of the right hand side of (4.10) for some auxiliary $a$, which will be the focus of the rest of the proof.
The structure of ${\rm G}_{{\mathfrak{c}},n}^{\prime}$. In this paragraph we analyse the structure of ${\rm G}_{{\mathfrak{c}},n}^{\prime}$ under our assumption that $\deg{\mathfrak{p}}=1$. A key result: ${\rm G}_{{\mathfrak{c}},n}^{\prime}$ only depends on $n$ and ${\rm G}_{{\mathfrak{c}},n}^{\prime}$ is independent of $n$.
Let $\Delta\subset{\rm{G}}_{{\mathfrak{c}}{\mathfrak{p}}^{\infty}}$ be the torsion subgroup. We have the short exact sequence
$0\rightarrow\Delta\rightarrow{\rm{G}}_{{\mathfrak{c}}{\mathfrak{p}}^{\infty}}\rightarrow\Gamma_{{\mathfrak{p}}}\rightarrow 0,$ (4.11)
where the third map is natural quotient map and $\Gamma_{\mathfrak{p}}\simeq{\mathbb{Z}}_{p}$. For $n>0$, let $\Gamma_{{\mathfrak{p}},n}$ be the quotient of $\Gamma_{\mathfrak{p}}$ by image of $1+{\mathfrak{p}}^{n}{\mathcal{O}}_{K_{\mathfrak{p}}}$ under the identification (4.5) and the quotient map in (4.11). Then for $n\gg 0$, (4.11) induces a short exact sequence
$0\rightarrow\Delta\rightarrow{\rm{G}}_{{\mathfrak{c}},n}\rightarrow\Gamma_{{\mathfrak{p}},n}\rightarrow 0,$ (4.12)
where the map ${\rm{G}}_{{\mathfrak{c}},n}\rightarrow\Gamma_{{\mathfrak{p}},n}$ is the natural quotient map. Since $\deg{\mathfrak{p}}=1$, we have
$\{\gamma\in\Gamma_{{\mathfrak{p}},n}\ |\ \nu(\gamma)\in\mu_{p^{s}}\subset\overline{{\mathbb{Q}}}\}=\Gamma_{{\mathfrak{p}},n}[p^{s}].$ (4.13)
Note that there exists $n_{0}$ such that for $n\geq n_{0}$, (4.12) holds and the natural map
$\frac{1+{\mathfrak{p}}^{n-s}O_{K_{\mathfrak{p}}}}{(1+{\mathfrak{p}}^{n}O_{K_{\mathfrak{p}}})(1+{\mathfrak{p}}^{n-s})}\rightarrow{\rm{P}}_{{\mathfrak{c}},n}\simeq{\rm{G}}_{{\mathfrak{c}},n}$ (4.14)
is injective, where the identification ${\rm{P}}_{{\mathfrak{c}},n}\simeq{\rm{G}}_{{\mathfrak{c}},n}$ is as in (4.6). So we have canonical identification
$\frac{1+{\mathfrak{p}}^{n-s}O_{K_{\mathfrak{p}}}}{(1+{\mathfrak{p}}^{n}O_{K_{\mathfrak{p}}})(1+{\mathfrak{p}}^{n-s})}\simeq\Gamma_{{\mathfrak{p}},n}[p^{s}]$ (4.15)
under the the map (4.14) and the quotient map in (4.12). Then it follows that
${\rm{G}}_{{\mathfrak{c}},n}^{\prime}=\Delta\times\frac{1+{\mathfrak{p}}^{n-s}O_{K_{\mathfrak{p}}}}{(1+{\mathfrak{p}}^{n}O_{K_{\mathfrak{p}}})(1+{\mathfrak{p}}^{n-s})}\subset{\rm G}_{{\mathfrak{c}},n}.$
Hence, ${\rm{G}}_{{\mathfrak{c}},n}^{\prime}$ depends only on $n$ and its cardinality is independent of $n$.
Also note that
$\Delta=D_{1}\times\Delta^{\text{alg}},$
where $D_{1}$, $\Delta^{\text{alg}}$ are as in §4.1.1. Define a subgroup
${\rm H}_{{\mathfrak{c}},n}:=\Delta^{{\mathrm{alg}}}\cdot\frac{1+{\mathfrak{p}}^{n-s}O_{K_{\mathfrak{p}}}}{(1+{\mathfrak{p}}^{n}O_{K_{\mathfrak{p}}})(1+{\mathfrak{p}}^{n-s})}\subset{\rm{G}}_{{\mathfrak{c}},n}^{\prime},$ (4.16)
which is the image of an open compact subgroup of $K_{S}^{\times}/F_{S}^{\times}$ for some finite set of places $S$ of $F$.
A Hecke-modification of the test vector outside $\mathfrak{p}$}. In this paragraph we introduce a Hecke-modification of the test vector at places dividing $\mathfrak{c}$}, which is eigen under the $\Delta^{\text{alg}}$-action.
Put $V_{0}^{({\mathfrak{p}})}=\prod_{v|{\mathfrak{c}}}O_{K_{v}}^{\times}/O_{K_{v},{\mathfrak{c}}_{v}}^{\times}$. Note that $V_{0}^{({\mathfrak{p}})}$ is an $\ell$-adic unit. Define
$H_{\ell}(g)=\frac{1}{\#V_{0}^{({\mathfrak{p}})}}\sum_{u\in V_{0}^{({\mathfrak{p}})}}\chi_{0}(u)\langle{x,\widehat{\varphi}(g\cdot\iota_{\varsigma_{\mathfrak{c}}^{({\mathfrak{c}})}}(u))}\rangle_{k-2\Sigma},\quad g\in\widehat{B}^{\times}.$ (4.17)
In view of the assumption $\ell\nmid\prod_{v|{\mathfrak{c}}}(q_{v}^{2}-1)$, Lemma 4.8 and [8, Prop. 3.12] [^22] Note that [8, Prop. 3.12] implies that the $\gamma$ in Lemma 4.8 is an $\ell$-adic unit since $\ell\nmid\prod_{v|{\mathfrak{c}}}q_{v}(q_{v}^{2}-1)$., it follows that $H_{\ell}(g)\pmod{\varpi}\in{\rm M}_{2}(U^{\prime},{\bf k}_{\ell})$ is also a non-zero CM form, where $U^{\prime}\subset\widehat{B}^{\times}$ is an open compact subgroup such that $U^{\prime({\mathfrak{c}})}=U^{({\mathfrak{c}})}$. By the definition of $\Delta^{{\mathrm{alg}}}$, we have
$\sum_{t\in{\rm G}_{{\mathfrak{c}},n}}\chi_{0}\nu(t)\langle{x,\widehat{\varphi}(x_{{\mathfrak{c}},n}(at))}\rangle_{k-2\Sigma}\equiv\sum_{t\in{\rm G}_{{\mathfrak{c}},n}}\chi_{0}\nu(t)H_{\ell}(x_{{\mathfrak{c}},n}(at))\pmod{\varpi},\qquad\forall a\in\widehat{K}^{\times}$ (4.18)
and
$\sum_{t\in{\rm G}_{{\mathfrak{c}},n}^{\prime}}\chi_{0}\nu(t)\langle{x,\widehat{\varphi}(x_{{\mathfrak{c}},n}(at))}\rangle_{k-2\Sigma}\equiv\sum_{t\in{\rm G}_{{\mathfrak{c}},n}^{\prime}}\chi_{0}\nu(t)H_{\ell}(x_{{\mathfrak{c}},n}(at))\pmod{\varpi},\qquad\forall a\in\widehat{K}^{\times}.$ (4.19)
In view of the definition of $f$, note that $H_{\ell}(x_{{\mathfrak{c}},n}(\cdot))$ is $\Delta^{\text{alg}}$-eigen with eigen-character $\chi_{0}^{-1}$, and $\Delta^{\text{alg}}$ is an $\ell$-adic unit by our assumption on $\ell$. Thus for $\nu\in\Xi_{\mathfrak{p}}$ with $n={\mathrm{ord}}_{\mathfrak{p}}\mathrm{cond}(\nu)\geq n_{0}$, we have
$\sum_{t\in{\rm G}_{{\mathfrak{c}},n}^{\prime}}\chi_{0}\nu(t)H_{\ell}(x_{{\mathfrak{c}},n}(at))\equiv\#\Delta^{\text{alg}}\sum_{t\in D_{1}\times\frac{1+{\mathfrak{p}}^{n-s}O_{K_{\mathfrak{p}}}}{(1+{\mathfrak{p}}^{n}O_{K_{\mathfrak{p}}})(1+{\mathfrak{p}}^{n-s})}}\chi_{0}\nu(t)H_{\ell}(x_{{\mathfrak{c}},n}(at))\pmod{\varpi},\qquad a\in\widehat{K}^{\times}.$ (4.20)
A partition of $\Xi_{{\mathfrak{p}}}$ and Hecke-modification of the test vector at $\mathfrak{p}$}. In this paragraph we partition $\Xi_{{\mathfrak{p}}}$ into finitely many subfamilies indexed by primitive additive characters $\mu:{\mathbb{Z}}/p^{s}{\mathbb{Z}}\twoheadrightarrow\mu_{p^{s}}\subset\overline{{\mathbb{Q}}}$, and introduce a Hecke-modification $H_{\ell,\mu}$ of $H_{\ell}$ at $\mathfrak{p}$} which is eigen under the action of $\frac{1+{\mathfrak{p}}^{n-s}O_{K_{\mathfrak{p}}}}{(1+{\mathfrak{p}}^{n}O_{K_{\mathfrak{p}}})(1+{\mathfrak{p}}^{n-s})}$.
We always identify $F_{{\mathfrak{p}}}$ with ${\mathbb{Q}}_{p}$ and the matrices appearing in the Hecke action are seen as elements of $B^{\times}(F_{\mathfrak{p}})\simeq{\mathrm{GL}}_{2}({\mathbb{Q}}_{p})$, where the isomorphism is as in §2.2. We choose a local uniformiser of $F_{\mathfrak{p}}$ to be $p$.
Let $\{1,\alpha\}$ be a relative integral basis of $O_{K_{\mathfrak{p}}}$ as in  §2.4. Note that there exists a sufficiently large integer $n_{1}\geq s$ such that for any integer $n\geq n_{1}$ the following holds: we have an isomorphism
${\mathbb{Z}}/p^{s}{\mathbb{Z}}\simeq\frac{1+{\mathfrak{p}}^{n-s}O_{K_{\mathfrak{p}}}}{(1+{\mathfrak{p}}^{n}O_{K_{\mathfrak{p}}})(1+{\mathfrak{p}}^{n-s})},\quad u\mapsto 1+\alpha p^{n-s}u.$
For any $u\in{\mathbb{Z}}/p^{s}{\mathbb{Z}}=O_{\mathfrak{p}}/{\mathfrak{p}}^{s}$, we have
$\iota(1+\alpha p^{n-s}u)\varsigma_{{\mathfrak{p}}}^{(n)}\equiv\varsigma_{{\mathfrak{p}}}^{(n)}\begin{pmatrix}1&\frac{u}{p^{s}}\\ &1\end{pmatrix}\pmod{U_{0}(\mathfrak{N}_{\lambda,{\mathfrak{p}}})_{\mathfrak{p}}}$ (4.21)
as elements in ${\mathrm{GL}}_{2}(F_{\mathfrak{p}})$. Put
$\nu(1+\alpha p^{n-s}u)=\zeta_{\nu}^{u}$
where
$\zeta_{\nu}:=\nu(1+\alpha p^{n-s})\in\mu_{p^{s}}\subset\overline{{\mathbb{Q}}}^{\times}$
is a primitive $p^{s}$-th root of unity.
In view of (4.21) and the definition of $\zeta_{\nu}$, it thus follows that
$\sum_{t\in\frac{1+p^{n-s}O_{K_{\mathfrak{p}}}}{(1+p^{n}O_{K_{\mathfrak{p}}})(1+{\mathfrak{p}}^{n-s})}}\nu(t)H_{\ell}(x_{{\mathfrak{c}},n}(at))=\sum_{y\in{\mathbb{Z}}/p^{s}\mathbb{Z}}\zeta_{\nu}^{y}H_{\ell}\left(x_{{\mathfrak{c}},n}(a\tau)\begin{pmatrix}1&\frac{y}{p^{s}}\\ 0&1\\ \end{pmatrix}\right)\qquad\forall a\in\widehat{K}^{\times}.$ (4.22)
Consequently, in view of (4.10), (4.19) and (4.20), for $n>m_{0}:=\max\{n_{0},n_{1}\}$ and $a\in\widehat{K}^{\times}$, we have
${\rm Tr}_{{\bf k}_{\ell}(\nu)/{\bf k}_{\ell}}(\nu\chi_{0}(a^{-1})\cdot\nu(\Theta_{\chi_{0},n}(\pi_{\lambda})\pmod{\varpi})$ (4.23)
$=$ $\#\Delta^{\text{alg}}d_{\nu}\cdot\sum_{\tau\in D_{1}}{\chi_{0}}(\tau)\sum_{y\in{\mathbb{Z}}/p^{s}\mathbb{Z}}\zeta_{\nu}^{y}H_{\ell}\left(x_{{\mathfrak{c}},n}(a\tau)\begin{pmatrix}1&\frac{y}{p^{s}}\\ 0&1\\ \end{pmatrix}\right)\pmod{\varpi})$.
Let
$\mu:{\mathbb{Z}}/p^{s}{\mathbb{Z}}\twoheadrightarrow\mu_{p^{s}}\subset\overline{{\mathbb{Q}}}^{\times}$
be a given character. Let $\Xi_{{\mathfrak{p}}}(\mu)\subset\Xi_{{\mathfrak{p}}}$ be the subset consisting of characters $\nu$ such that ${\mathrm{ord}}_{\mathfrak{p}}\mathrm{cond}(\nu)>m_{0}$ and $\zeta_{\nu}^{\cdot}=\mu(\cdot)$. Define
$\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{\pm}(\mu)=\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{\pm}\cap\Xi_{{\mathfrak{p}}}(\mu).$
Note that $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}(\mu)$ and $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{-}(\mu)$ have infinitely elements by Lemma 4.12 since $\mathfrak{p}$} is inert.
Define
$H_{\ell,\mu}(g)=\frac{1}{p^{s}}\sum_{y\in{\mathbb{Z}}/p^{s}{\mathbb{Z}}}\mu(y)\pi_{\lambda,{\mathfrak{p}}}\left(\begin{pmatrix}1&\frac{y}{p^{s}}\\ 0&1\\ \end{pmatrix}\right)H_{\ell}(g),\quad g\in\widehat{B}^{\times}.$ (4.24)
Note that for each $\nu\in\Xi_{\mathfrak{p}}(\mu)$ and $n:={\mathrm{ord}}_{\mathfrak{p}}\mathrm{cond}(\nu)$, the form $H_{\ell,\mu}(x_{{\mathfrak{c}},n}(\cdot))$ is $(\chi_{0}\nu)^{-1}$-eigen under the action of ${\rm H}_{{\mathfrak{c}},n}$ as in (4.16). In view of the eigen-property, the remark just below (4.16) and Lemma 3.12, it follows that
$H_{\ell,\mu}\neq 0.$
By definition, we have $H_{\ell,\mu}\pmod{\varpi}\in{\rm M}_{2}(\widetilde{U}^{\prime},{\bf k}_{\ell}(\nu))$ for $\widetilde{U}^{\prime}:=U^{\prime({\mathfrak{p}})}\cdot U_{1}({\mathfrak{p}}^{k_{0}})$. Let $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}(\mu)$. Then in view of (4.23), we have
${\rm Tr}_{{\bf k}_{\ell}(\nu)/{\bf k}_{\ell}}(\nu\chi_{0}(a^{-1})\cdot\nu(\Theta_{\chi_{0},n}(\pi_{\lambda}))\equiv\#\Delta^{{\mathrm{alg}}}p^{s}d_{\nu}\cdot\sum_{\tau\in D_{1}}{\chi_{0}}(\tau)H_{\ell,\mu}(x_{{\mathfrak{c}},n}(a\tau))\pmod{\varpi}$. (4.25)
Being non-Eisenstein. We next show that $H_{\ell,\mu}\pmod{\varpi}$ is non-Eisenstein.
To begin, (4.4) holds by Lemma 3.11. Therefore, since $\pi_{\lambda}$ is CM and $(\ell,2D_{K/F})=1$, the non-vanishing of $H_{\ell}\pmod{\varpi}$ implies that it is non-Eisenstein by Lemma 4.10. A simple calculation shows that
$\sum_{t\in(\mathbb{Z}/p^{s}\mathbb{Z})^{\times}}\rho\left(\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\right)H_{\ell,\mu}=$ $p^{s}H_{\ell}-{p^{s-1}}\rho\left(\begin{pmatrix}p^{-1}&\\ &1\end{pmatrix}\right)U_{\mathfrak{p}}H_{\ell}.$ (4.26)
$\sum_{t\in(\mathbb{Z}/p^{s}\mathbb{Z})^{\times}}\rho\left(\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\right)H_{\ell,\mu}=p^{s}H_{\ell}-{p^{s-1}}\rho\left(\begin{pmatrix}p^{-1}&\\ &1\end{pmatrix}\right)T_{\mathfrak{p}}H_{\ell}+{p^{s-1}}\rho\left(\begin{pmatrix}p^{-2}&\\ &1\end{pmatrix}\right)H_{\ell}.$ (4.27)
Since $H_{\ell}$ is either a $U_{\mathfrak{p}}$-eigenform or a $T_{\mathfrak{p}}$-eigenform with an $\ell$-integral eigenvalue and $H_{\ell}\pmod{\varpi}$ is non-Eisenstein, it follows that $H_{\ell,\mu}\pmod{\varpi}$ is non-Eisenstein by Lemma 4.7. Then we may suppose that $H_{\ell,\mu}^{\varepsilon}\pmod{\varpi}$ is non-Eisenstein for some $\varepsilon\in\{+,-\}$.
The non-vanishing. In view of (4.25) and Proposition 4.5, it thus follows that
@@
for all but finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{{\varepsilon_{0}}}(\mu)$ such that $n:={\mathrm{ord}}_{\mathfrak{p}}\mathrm{cond}(\nu)$ satisfies
$\varepsilon=(-1)^{n}\eta_{K/F}({\mathfrak{c}}).$ (4.28)
Note that $\varepsilon_{0}:=\varepsilon(\lambda\chi_{0}\nu)=(-1)^{n-m_{\lambda}}\eta_{K/F}({\mathfrak{c}})\varepsilon(\lambda)$ by Lemma 4.12. So the condition (4.28) is equivalent to
$\varepsilon_{0}=(-1)^{m_{\lambda}}\varepsilon(\lambda){\varepsilon}.$ (4.29)
@@
for all but finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{-}(\mu)$. But these central $L$-values vanish since $\varepsilon(\lambda\chi_{0}\nu)=-1$ for any $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{-}(\mu)$ and $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{-}(\mu)=\infty$. Therefore $\varepsilon_{0}=+$ and the parity $n$ in Proposition 4.5 satisfies $(-1)^{n-m_{\lambda}}\varepsilon(\lambda\chi_{0})=+1$. Moreover, it follows from (4.29) that ${H}_{\ell,\mu}^{{\mathrm{sgn}}((-1)^{m_{\lambda}}\varepsilon(\lambda))}\pmod{\varpi}$ is non-Eisenstein and ${H}_{\ell,\mu}^{{\mathrm{sgn}}(-(-1)^{m_{\lambda}}\varepsilon(\lambda))}\pmod{\varpi}$ is Eisenstein, where ${\mathrm{sgn}}(\pm 1):=\pm$. ∎
Remark 4.16.
If $\deg{\mathfrak{p}}>1$, then the group ${\rm G}_{{\mathfrak{c}},n}^{\prime}$ appearing in (4.10) is more complex: it has a significant contribution at $\mathfrak{p}$} given by $\ker(\nu)$ since $\Gamma_{n,{\mathfrak{p}}}$ is a direct sum of $\deg{\mathfrak{p}}$-many cyclic $p$-groups. In particular ${\rm G}_{{\mathfrak{c}},n}^{\prime}$ also depends on $n$. Consequently, the existence of an analogous common test vector $H_{\ell,\mu}$ as in the above proof is not apparent, and it will be considered in the sequel [4].